### v1

@@ 7 Conclusion and future outlook
The motivation for this section is twofold. First, it generalizes our previous work from [16], which used MGRIT to solve method-of-lines discretizations of constant wave-speed advection problems, because we now solve variable wave-speed conservation laws. Second, and most relevant to the main paper, our development in the next section, Section SM6, of the truncation error operator ${\cal T}_{\rm ideal}^{n}$ used in (58) for the MGRIT solution of linearized method-of-lines discretizations is heavily inspired by the development of the ${\cal T}_{\rm ideal}^{n}$ in this section for standard, linear method-of-lines discretizations.
The remainder of this section is organized as follows. Section SM5.1 gives details of the fine-grid discretization and presents error estimates required to approximate the truncation error. Section SM5.2 presents the coarse-grid operator, then numerical MGRIT results follow in Section SM5.3.
SM5.1 Fine-grid discretization and required error estimates
@@
    $\widehat{f}_{i+1/2}$ $=\frac{1}{2}\left[\alpha(x_{i+1/2})\big{(}e_{i+1/2}^{-}+e_{i+1/2}^{+}\big{)}+\nu_{i+1/2}\big{(}e_{i+1/2}^{-}-e_{i+1/2}^{+}\big{)}\right],$ (SM54)
    $=\alpha(x_{i+1/2})e(x_{i+1/2})+\widehat{E}_{i+1/2},$ (SM55)
@@
  • The reconstructions $e_{i+1/2}^{\pm}$ in the numerical flux (SM54) are computed with polynomial reconstruction on central stencils using $12k-$ 1 cell averages (specifics of polynomial reconstructions can be recalled from Sections 2.2 and SM1). Thus, for sufficiently smooth $e(x)$, the reconstructions are of order $p=2k-1$, $k=1,2,\ldots$.
  • If using GLF, the dissipation coefficient in (SM54) is $\nu^{\rm(global)}=\max_{x,t}|\alpha(x,t)|$, with the range of $x,t$ taken as the whole space-time domain, and if using LLF $\nu_{i+1/2}^{n}=|\alpha(x_{i+1/2},t_{n})|$.
With details of the discretization specified, we now estimate the error term in (SM55) such that the analysis from Section SM4 can be applied to estimate the truncation error of the ideal coarse-grid operator. To this end, consider the following corollary of Lemma SM4.5 to express the reconstructions used in (SM54).
Corollary SM5.1 (Error estimate for central, linear reconstructions using $12k-$ 1 cells).
Let $e_{i,\pm 1/2}$ be polynomial reconstructions of the function $e(x_{i\pm 1/2})$ using central stencils containing $12k-$ 1 cells, $S(i)=\{{\cal I}_{i-(k-1)},\ldots,{\cal I}_{i+(k-1)}\}$. Suppose that $e(x)$ is $2k$ times differentiable over the cells in this stencil. Then,
$e_{i,-1/2}$ $=e(x_{i-1/2})-(-1)^{k}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i}(x_{i-1/2})},$ (SM56)
$e_{i,+1/2}$ $=e(x_{i+1/2})-(-1)^{k+1}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i}(x_{i+1/2})},$ (SM57)
@@
Proof SM5.2.
The central stencil considered here corresponds to the stencil from Lemma SM4.5 with a number cells $k\mapsto 2k-1$, and left shift $\ell=k-1$. Plugging these parameters into the results of Lemma SM4.5 and simplifying gives the claim.
Now we estimate the error term of the numerical flux, as it appears in (SM55).
Lemma SM5.3 (Error estimate for numerical flux using central linear reconstructions).
Consider the numerical flux (SM54), and suppose that the polynomial reconstructions $e_{i+1/2}^{\pm}$ use central stencils containing $12k-$ 1 cells. Suppose that $e(x)$ is sufficiently smooth over the stencils in question. Then, the error of the numerical flux, as defined in (SM55), satisfies
$\widehat{E}_{i+1/2}=\frac{\nu_{i+1/2}}{2}\big{(}e_{i+1/2}^{-}-e_{i+1/2}^{+}\big{)}+{\cal O}(h^{2k}),$ (SM58)
@@
$e_{i+1/2}^{-}-e_{i+1/2}^{+}=(-1)^{k}h^{2k-1}\frac{2(k-1)!k!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{x_{i+1/2}}+{\cal O}(h^{2k}).$ (SM59)
Proof SM5.4.
Using the notation of Corollary SM5.1, the reconstructions considered in the lemma are $e_{i+1/2}^{-}=e_{i,+1/2}$, and $e_{i+1/2}^{+}=e_{i+1,-1/2}$. First consider the sum $e_{i+1/2}^{-}+e_{i+1/2}^{+}$ in the flux (SM54); plugging in the estimates from Corollary SM5.1 we have
$\begin{split}&e_{i+1/2}^{-}+e_{i+1/2}^{+}=2e(x_{i+1/2})\\ &\quad+(-1)^{k+1}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\left[\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i+1}(x_{i+1/2})}-\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i}(x_{i+1/2})}\right],\end{split}$ (SM60)
@@
$\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i+1}(x_{i+1/2})}-\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i}(x_{i+1/2})}=\frac{\operatorname{d}\!^{2k}e}{\operatorname{d}\!x^{2k}}\bigg{|}_{c}\big{[}\xi_{i+1}(x_{i+1/2})-\xi_{i}(x_{i+1/2})\big{]}$ (SM61)
for some $\min\big{(}\xi_{i+1}(x_{i+1/2}),\xi_{i}(x_{i+1/2})\big{)}<c<\max\big{(}\xi_{i+1}(x_{i+1/2}),\xi_{i}(x_{i+1/2})\big{)}$. Next, note that $\xi_{i+1}(x_{i+1/2})-\xi_{i}(x_{i+1/2})={\cal O}(h)$ because both of these points are contained within an interval of length ${\cal O}(h)$. Plugging into (SM60) gives $e_{i+1/2}^{-}+e_{i+1/2}^{+}=2e(x_{i+1/2})+{\cal O}(h^{2k})$.
Now consider the difference $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ in the numerical flux (SM54). Using the estimates from Corollary SM5.1 we find
$e_{i+1/2}^{-}-e_{i+1/2}^{+}$ $=(-1)^{k}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\left[\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i+1}(x_{i+1/2})}+\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{\xi_{i}(x_{i+1/2})}\right],$ (SM62)
$=(-1)^{k}h^{2k-1}\frac{2(k-1)!k!}{(2k)!}\left[\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{x_{i+1/2}}+{\cal O}(h)\right],$ (SM63)
with the second expression following by Taylor expansion, again, due to the fact that $\xi_{i+1}(x_{i+1/2})+{\cal O}(h)=\xi_{i}(x_{i+1/2})+{\cal O}(h)=x_{i+1/2}$. Plugging (SM63) and the earlier result of $e_{i+1/2}^{-}+e_{i+1/2}^{+}=2e(x_{i+1/2})+{\cal O}(h^{2k})$ into (SM54) concludes the proof.
SM5.2 Modified semi-Lagrangian coarse-grid operator
@@
$\Psi^{n}=\Big{[}I+{\cal T}_{\rm ideal}^{n}-{\cal T}_{\rm direct}^{n}\Big{]}^{-1}{\cal S}^{n,m\delta t}_{p,1_{*}}\approx\Psi_{\rm ideal}^{n}.$ (SM64)
The approximate, direct truncation operator ${\cal T}_{\rm direct}^{n}$ is given by (SM28)—more generally, see Section SM3.3 for details on how this approximation arises. We now discuss our choice for the approximate the ideal truncation error operator ${\cal T}_{\rm ideal}^{n}$. Other details on the coarse-grid operator (SM64) can be found in Section 5.2.
Recall Eq. SM46, which is an approximate error estimate for a method-of-lines discretization for an approximation to the PDE at hand (cons-lin) in which the wave-speed function $\alpha(x,t)$ is frozen somewhere inside each of the time intervals $t\in[t_{n+j},t_{n+1+j}]$, $j\in\{0,\ldots,m-1\}$. Based on this error estimate, which applies approximately to the problem at hand, we choose ${\cal T}_{\rm ideal}^{n}$ to take the form
${\cal T}_{\rm ideal}^{n}=\textrm{approximation}\bigg{(}\sum_{j=0}^{m-1}\Big{(}\delta t\widehat{{\cal E}}^{n+j}-\widehat{e}_{\rm RK}\big{(}\delta t\widehat{L}^{n+j}\big{)}^{q+1}\Big{)}\bigg{)}.$ (SM65)
The “approximation” here represents the fact that the terms in the sum need to be approximated, and that the choice for doing so is not unique. Recall that $\widehat{L}^{n+j}\in\mathbb{R}^{n_{x}\times n_{x}}$ is the spatial discretization at some time $t=t_{n+j}+\delta t\vartheta$, and ^$\widehat{{\cal E}}^{n+j}\in\mathbb{R}^{n_{x}\times n_{x}}$ is the associated spatial discretization error at this time; see Footnote 10. The question at hand is how to approximate $\widehat{L}^{n+j}$ and ^$\widehat{{\cal E}}^{n+j}$. It is perhaps simplest to start with ^$\widehat{{\cal E}}^{n+j}$.
Approximating the spatial discretization error. The exact form of ^$\widehat{{\cal E}}^{n+j}$ depends on the numerical flux. Recall from Lemma SM5.3 of Section SM5.1 that the numerical flux at hand satisfies the estimate (dropping temporal superscripts for readability)
$\widehat{f}_{i+1/2}=\alpha(x_{i+1/2})e(x_{i+1/2})+\nu_{i+1/2}(-1)^{k}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{x_{i+1/2}}+{\cal O}(h^{2k}).$ (SM66)
@@
(^$\big{(}\widehat{{\cal E}}\bar{\bm{e}}\big{)}_{i}$ $=-\frac{\alpha(x_{i+1/2})e(x_{i+1/2})-\alpha(x_{i-1/2})e(x_{i-1/2})}{h}+\frac{\widehat{f}_{i+1/2}-\widehat{f}_{i-1/2}}{h},$ (SM67)
$\approx(-1)^{k+1}h^{2k-2}\frac{k!(k-1)!}{(2k)!}\left[\nu_{i+1/2}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{x_{i+1/2}}-\nu_{i-1/2}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg{|}_{x_{i-1/2}}\right].$ (SM68)
Whether or not the terms enumerated here accurately capture the largest terms in the (^$\big{(}\widehat{{\cal E}}\bar{\bm{e}}\big{)}_{i}$ depends on the smoothness of the numerical flux. [^11] That is, without further analysis, it is not clear whether the difference of the ${\cal O}(h^{2k})$ terms present in both $\widehat{f}_{i+1/2}$ and $\widehat{f}_{i-1/2}$ leads to terms which are of the same size as those enumerated here, or whether they are one order smaller. In any event, we find this expression is sufficient for our purposes. We then implement numerically an approximation to this formula derived from approximating the degree $2k-1=p$ derivatives with finite differences. Specifically, we use
$\sum_{j=0}^{m-1}\widehat{{\cal E}}^{n+j}\approx(-1)^{k+1}h^{p}\frac{k!(k-1)!}{(2k)!}{\cal D}_{1}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\bigg{)}{\cal D}_{p}^{\top},$ (SM69)
where $\bm{\nu}^{n+j}=\big{(}\nu_{1+1/2}^{n+j},\nu_{2+1/2}^{n+j},\ldots,\nu_{n_{x}+1/2}^{n+j}\big{)}\in\mathbb{R}^{n_{x}}$ is the vector of dissipation coefficients used in the LF numerical flux (12) at time $t_{n+j}+\delta t\vartheta$. Recall from Section SM4.1 that to develop a truncation error estimate for the linear problem, its time dependence is frozen somewhere in the interval $[t_{n},t_{n}+\delta t]$, and the parameter $\vartheta\in[0,1]$ controls the specific freezing point.
Approximating powers of the spatial discretization. It is less obvious how to approximate $\big{(}\widehat{L}^{n+j}\big{)}^{q+1}$, the term in (SM65) corresponding to the ($q+1$)st power of the spatial discretization matrix. We do not want to compute this exactly, since the implementation of the discretization does not even require computing $\widehat{L}^{n+j}$ as a matrix, so, explicitly forming it and then raising it to the ($q+1$)st power seems too expensive. Thus, we seek to cheaply approximate it in a way that is consistent with the approximation that we have for the spatial error matrix in (SM69).
@@
$\big{(}\widehat{L}\bar{\bm{e}}\big{)}_{i}$ $=-\frac{\alpha(x_{i+1/2})e(x_{i+1/2})-\alpha(x_{i-1/2})e(x_{i-1/2})}{h}+\rm{h.o.t}.,$ (SM70)
$=-\frac{\partial}{\partial x}\Big{(}\alpha(x)e(x)\Big{)}\bigg{|}_{x_{i}}+\rm{h.o.t}.$ (SM71)
@@
$\Big{(}(\widehat{L})^{q+1}\bar{\bm{e}}\Big{)}_{i}=(-1)^{q+1}\underbrace{\frac{\partial}{\partial x}\Big{(}\alpha(x)\Big{(}\cdots\frac{\partial}{\partial x}\Big{(}\alpha(x)}_{q\textrm{ terms}}\Big{(}\frac{\partial}{\partial x}\Big{(}\alpha(x)e(x)\Big{)}\Big{)}\Big{)}\cdots\Big{)}\Big{)}\bigg{|}_{x_{i}}+\rm{h.o.t}.$ (SM72)
@@
$\Big{(}\widehat{L}^{q+1}\bar{\bm{e}}\Big{)}_{i}\approx(-1)^{q+1}\frac{\partial}{\partial x}\Big{(}\alpha^{q+1}\frac{\partial^{q}e}{\partial x^{q}}\Big{)}\bigg{|}_{x_{i}}=(-1)^{q+1}\Big{(}{\cal D}_{1}\operatorname{diag}(\bm{\alpha})^{q+1}{\cal D}_{q}^{\top}\bar{\bm{e}}\Big{)}_{i}+\rm{h.o.t}.$ (SM73)
In this final expression, $\bm{\alpha}=\big{(}\alpha_{1+1/2},\ldots,\alpha_{n_{x}+1/2}\big{)}^{\top}\in\mathbb{R}^{n_{x}}$ is a vector comprised of the wave-speed function evaluated at cell interfaces and powers of it are computed element-wise. Thus, for the second term in (SM65) we use the approximation
$\sum\limits_{j=0}^{m-1}\big{(}\widehat{L}^{n+j}\big{)}^{q+1}\approx(-1)^{q+1}{\cal D}_{1}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\big{(}\bm{\alpha}^{n+j}\big{)}^{q+1}\bigg{)}{\cal D}_{q}^{\top},$ (SM74)
@@
Piecing it all together. Plugging (SM69) and (SM74) into (SM65) results in the form of ${\cal T}_{\rm ideal}^{n}$ that we implement numerically:
$\begin{split}{\cal T}_{\rm ideal}^{n}&={\cal D}_{1}\bigg{[}h^{p}\delta t(-1)^{k+1}\frac{k!(k-1)!}{(2k)!}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\bigg{)}{\cal D}_{p}^{\top}\\ &\hskip 51.6665pt+(-\delta t)^{q+1}\widehat{e}_{\rm RK}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\big{(}\bm{\alpha}^{n+j}\big{)}^{q+1}\bigg{)}{\cal D}_{q}^{\top}\bigg{]}.\end{split}$ (SM75)
SM5.3 Numerical results

### v2

@@ Acknowledgments
The motivation for this section is twofold. First, it generalizes our previous work from [15], which used MGRIT to solve method-of-lines discretizations of constant wave-speed advection problems, because we now solve variable wave-speed conservation laws. Second, and most relevant to the main paper, our development in the next section, Section SM8, of the truncation error operator ${\cal T}_{\rm ideal}^{n}$ used in (40) for the MGRIT solution of linearized method-of-lines discretizations is heavily inspired by the development of the ${\cal T}_{\rm ideal}^{n}$ in this section for standard, linear method-of-lines discretizations.
The remainder of this section is organized as follows. Section SM7.1 gives details of the fine-grid discretization and presents error estimates required to approximate the truncation error. Section SM7.2 presents the coarse-grid operator, then numerical MGRIT results follow in Section SM7.3.
SM7.1 Fine-grid discretization and required error estimates
@@
    (SM64) $\widehat{f}_{i+1/2}$ $=\frac{1}{2}\left[\alpha(x_{i+1/2})\big(e_{i+1/2}^{-}+e_{i+1/2}^{+}\big)+\nu_{i+1/2}\big(e_{i+1/2}^{-}-e_{i+1/2}^{+}\big)\right],$
    (SM65) $=\alpha(x_{i+1/2})e(x_{i+1/2})+\widehat{E}_{i+1/2},$
@@
  • The reconstructions $e_{i+1/2}^{\pm}$ in the numerical flux (SM64) are computed with polynomial reconstruction on central stencils using 2k12k-1 cell averages (specifics of polynomial reconstructions can be recalled from Sections 2.2 and SM1). Thus, for sufficiently smooth $e(x)$, the reconstructions are of order p=2k1p=2k-1, $k=1,2,\ldots$.
  • If using GLF, the dissipation coefficient in (SM64) is $\nu^{\rm(global)}=\max_{x,t}|\alpha(x,t)|$, with the range of $x,t$ taken as the whole space-time domain, and if using LLF $\nu_{i+1/2}^{n}=|\alpha(x_{i+1/2},t_{n})|$.
With details of the discretization specified, we now estimate the error term in (SM65) such that the analysis from Section SM6 can be applied to estimate the truncation error of the ideal coarse-grid operator. To this end, consider the following corollary of Lemma SM6.5 to express the reconstructions used in (SM64).
Corollary SM7.1 (Error estimate for central, linear reconstructions using 2k12k-1 cells).
Let $e_{i,\pm 1/2}$ be polynomial reconstructions of the function $e(x_{i\pm 1/2})$ using central stencils containing 2k12k-1 cells, $S(i)=\{{\cal I}_{i-(k-1)},\ldots,{\cal I}_{i+(k-1)}\}$. Suppose that $e(x)$ is $2k$ times differentiable over the cells in this stencil. Then,
(SM66) $e_{i,-1/2}$ $=e(x_{i-1/2})-(-1)^{k}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i}(x_{i-1/2})},$
(SM67) $e_{i,+1/2}$ $=e(x_{i+1/2})-(-1)^{k+1}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i}(x_{i+1/2})},$
@@
Proof SM7.2.
The central stencil considered here corresponds to the stencil from Lemma SM6.5 with a number cells $k\mapsto 2k-1$, and left shift $\ell=k-1$. Plugging these parameters into the results of Lemma SM6.5 and simplifying gives the claim.
Now we estimate the error term of the numerical flux, as it appears in (SM65).
Lemma SM7.3 (Error estimate for numerical flux using central linear reconstructions).
Consider the numerical flux (SM64), and suppose that the polynomial reconstructions $e_{i+1/2}^{\pm}$ use central stencils containing 2k12k-1 cells. Suppose that $e(x)$ is sufficiently smooth over the stencils in question. Then, the error of the numerical flux, as defined in (SM65), satisfies
(SM68) $\widehat{E}_{i+1/2}=\frac{\nu_{i+1/2}}{2}\big(e_{i+1/2}^{-}-e_{i+1/2}^{+}\big)+{\cal O}(h^{2k}),$
@@
(SM69) $e_{i+1/2}^{-}-e_{i+1/2}^{+}=(-1)^{k}h^{2k-1}\frac{2(k-1)!k!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{x_{i+1/2}}+{\cal O}(h^{2k}).$
Proof SM7.4.
Using the notation of Corollary SM7.1, the reconstructions considered in the lemma are $e_{i+1/2}^{-}=e_{i,+1/2}$, and $e_{i+1/2}^{+}=e_{i+1,-1/2}$. First consider the sum $e_{i+1/2}^{-}+e_{i+1/2}^{+}$ in the flux (SM64); plugging in the estimates from Corollary SM7.1 we have
(SM70) $\begin{split}&e_{i+1/2}^{-}+e_{i+1/2}^{+}=2e(x_{i+1/2})\\ &\quad+(-1)^{k+1}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\left[\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i+1}(x_{i+1/2})}-\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i}(x_{i+1/2})}\right],\end{split}$
@@
(SM71) $\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i+1}(x_{i+1/2})}-\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i}(x_{i+1/2})}=\frac{\operatorname{d}\!^{2k}e}{\operatorname{d}\!x^{2k}}\bigg|_{c}\big[\xi_{i+1}(x_{i+1/2})-\xi_{i}(x_{i+1/2})\big]$
for some $\min\big(\xi_{i+1}(x_{i+1/2}),\xi_{i}(x_{i+1/2})\big)<c<\max\big(\xi_{i+1}(x_{i+1/2}),\xi_{i}(x_{i+1/2})\big)$. Next, note that $\xi_{i+1}(x_{i+1/2})-\xi_{i}(x_{i+1/2})={\cal O}(h)$ because both of these points are contained within an interval of length ${\cal O}(h)$. Plugging into (SM70) gives $e_{i+1/2}^{-}+e_{i+1/2}^{+}=2e(x_{i+1/2})+{\cal O}(h^{2k})$.
Now consider the difference $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ in the numerical flux (SM64). Using the estimates from Corollary SM7.1 we find
(SM72) $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ $=(-1)^{k}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\left[\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i+1}(x_{i+1/2})}+\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{\xi_{i}(x_{i+1/2})}\right],$
(SM73) $=(-1)^{k}h^{2k-1}\frac{2(k-1)!k!}{(2k)!}\left[\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{x_{i+1/2}}+{\cal O}(h)\right],$
with the second expression following by Taylor expansion, again, due to the fact that $\xi_{i+1}(x_{i+1/2})+{\cal O}(h)=\xi_{i}(x_{i+1/2})+{\cal O}(h)=x_{i+1/2}$. Plugging (SM73) and the earlier result of $e_{i+1/2}^{-}+e_{i+1/2}^{+}=2e(x_{i+1/2})+{\cal O}(h^{2k})$ into (SM64) concludes the proof.
SM7.2 Modified semi-Lagrangian coarse-grid operator
@@
(SM74) $\Psi^{n}=\Big[I+{\cal T}_{\rm ideal}^{n}-{\cal T}_{\rm direct}^{n}\Big]^{-1}{\cal S}^{n,m\delta t}_{p,1_{*}}\approx\Psi_{\rm ideal}^{n}.$
The approximate, direct truncation operator ${\cal T}_{\rm direct}^{n}$ is given by (SM40)—more generally, see Section SM5.3 for details on how this approximation arises. We now discuss our choice for the approximate the ideal truncation error operator ${\cal T}_{\rm ideal}^{n}$. Other details on the coarse-grid operator (SM74) can be found in Section 5.2.
Recall Eq. SM58, which is an approximate error estimate for a method-of-lines discretization for an approximation to the PDE at hand (cons-lin) in which the wave-speed function $\alpha(x,t)$ is frozen somewhere inside each of the time intervals $t\in[t_{n+j},t_{n+1+j}]$, $j\in\{0,\ldots,m-1\}$. Based on this error estimate, which applies approximately to the problem at hand, we choose ${\cal T}_{\rm ideal}^{n}$ to take the form
(SM75) ${\cal T}_{\rm ideal}^{n}=\textrm{approximation}\bigg(\sum_{j=0}^{m-1}\Big(\delta t\widehat{{\cal E}}^{n+j}-\widehat{e}_{\rm RK}\big(\delta t\widehat{L}^{n+j}\big)^{q+1}\Big)\bigg).$
The “approximation” here represents the fact that the terms in the sum need to be approximated, and that the choice for doing so is not unique. Recall that $\widehat{L}^{n+j}\in\mathbb{R}^{n_{x}\times n_{x}}$ is the spatial discretization at some time $t=t_{n+j}+\delta t\vartheta$, and $\widehat{{\cal E}}^{n+j}\in\mathbb{R}^{n_{x}\times n_{x}}$ is the associated spatial discretization error at this time; see Footnote 10. The question at hand is how to approximate $\widehat{L}^{n+j}$ and $\widehat{{\cal E}}^{n+j}$. It is perhaps simplest to start with $\widehat{{\cal E}}^{n+j}$.
Approximating the spatial discretization error. The exact form of $\widehat{{\cal E}}^{n+j}$ depends on the numerical flux. Recall from Lemma SM7.3 of Section SM7.1 that the numerical flux at hand satisfies the estimate (dropping temporal superscripts for readability)
(SM76) $\widehat{f}_{i+1/2}=\alpha(x_{i+1/2})e(x_{i+1/2})+\nu_{i+1/2}(-1)^{k}h^{2k-1}\frac{k!(k-1)!}{(2k)!}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{x_{i+1/2}}+{\cal O}(h^{2k}).$
@@
(SM77) $\big(\widehat{{\cal E}}\bar{\bm{e}}\big)_{i}$ $=-\frac{\alpha(x_{i+1/2})e(x_{i+1/2})-\alpha(x_{i-1/2})e(x_{i-1/2})}{h}+\frac{\widehat{f}_{i+1/2}-\widehat{f}_{i-1/2}}{h},$
(SM78) $\approx(-1)^{k+1}h^{2k-2}\frac{k!(k-1)!}{(2k)!}\left[\nu_{i+1/2}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{x_{i+1/2}}-\nu_{i-1/2}\frac{\operatorname{d}\!^{2k-1}e}{\operatorname{d}\!x^{2k-1}}\bigg|_{x_{i-1/2}}\right].$
Whether or not the terms enumerated here accurately capture the leading-order terms in the $\big(\widehat{{\cal E}}\bar{\bm{e}}\big)_{i}$ depends on the smoothness of the numerical flux. [^11] That is, without further analysis, it is not clear whether the difference of the ${\cal O}(h^{2k})$ terms present in both $\widehat{f}_{i+1/2}$ and $\widehat{f}_{i-1/2}$ leads to terms which are of the same size as those enumerated here, or whether they are one order smaller. In any event, we find this expression is sufficient for our purposes. We then implement numerically an approximation to this formula derived from approximating the degree 2k1=p2k-1=p derivatives with finite differences. Specifically, we use
(SM79) $\sum_{j=0}^{m-1}\widehat{{\cal E}}^{n+j}\approx(-1)^{k+1}h^{p}\frac{k!(k-1)!}{(2k)!}{\cal D}_{1}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\bigg){\cal D}_{p}^{\top},$
where $\bm{\nu}^{n+j}=\big(\nu_{1+1/2}^{n+j},\nu_{2+1/2}^{n+j},\ldots,\nu_{n_{x}+1/2}^{n+j}\big)\in\mathbb{R}^{n_{x}}$ is the vector of dissipation coefficients used in the LF numerical flux (10) at time $t_{n+j}+\delta t\vartheta$. Recall from Section SM6.1 that to develop a truncation error estimate for the linear problem, its time dependence is frozen somewhere in the interval $t_{n},t_{n}+\delta t]$, and the parameter $\vartheta\in[0,1]$ controls the specific freezing point.
Approximating powers of the spatial discretization. It is less obvious how to approximate $\big(\widehat{L}^{n+j}\big)^{q+1}$, the term in (SM75) corresponding to the ($q+1$)st power of the spatial discretization matrix. We do not want to compute this exactly, since the implementation of the discretization does not even require computing $\widehat{L}^{n+j}$ as a matrix, so, explicitly forming it and then raising it to the ($q+1$)st power seems too expensive. Thus, we seek to cheaply approximate it in a way that is consistent with the approximation that we have for the spatial error matrix in (SM79).
@@
(SM80) $\big(\widehat{L}\bar{\bm{e}}\big)_{i}$ $=-\frac{\alpha(x_{i+1/2})e(x_{i+1/2})-\alpha(x_{i-1/2})e(x_{i-1/2})}{h}+\rm{h.o.t}.,$
(SM81) $=-\frac{\partial}{\partial x}\Big(\alpha(x)e(x)\Big)\bigg|_{x_{i}}+\rm{h.o.t}.$
@@
(SM82) $\Big((\widehat{L})^{q+1}\bar{\bm{e}}\Big)_{i}=(-1)^{q+1}\underbrace{\frac{\partial}{\partial x}\Big(\alpha(x)\Big(\cdots\frac{\partial}{\partial x}\Big(\alpha(x)}_{q\textrm{ terms}}\Big(\frac{\partial}{\partial x}\Big(\alpha(x)e(x)\Big)\Big)\Big)\cdots\Big)\Big)\bigg|_{x_{i}}+\rm{h.o.t}.$
@@
(SM83) $\Big(\widehat{L}^{q+1}\bar{\bm{e}}\Big)_{i}\approx(-1)^{q+1}\frac{\partial}{\partial x}\Big(\alpha^{q+1}\frac{\partial^{q}e}{\partial x^{q}}\Big)\bigg|_{x_{i}}=(-1)^{q+1}\Big({\cal D}_{1}\operatorname{diag}(\bm{\alpha})^{q+1}{\cal D}_{q}^{\top}\bar{\bm{e}}\Big)_{i}+\rm{h.o.t}.$
In this final expression, $\bm{\alpha}=\big(\alpha_{1+1/2},\ldots,\alpha_{n_{x}+1/2}\big)^{\top}\in\mathbb{R}^{n_{x}}$ is a vector comprised of the wave-speed function evaluated at cell interfaces and powers of it are computed element-wise. Thus, for the second term in (SM75) we use the approximation
(SM84) $\sum\limits_{j=0}^{m-1}\big(\widehat{L}^{n+j}\big)^{q+1}\approx(-1)^{q+1}{\cal D}_{1}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\big(\bm{\alpha}^{n+j}\big)^{q+1}\bigg){\cal D}_{q}^{\top},$
@@
Piecing it all together. Plugging (SM79) and (SM84) into (SM75) results in the form of ${\cal T}_{\rm ideal}^{n}$ that we implement numerically:
(SM85) $\begin{split}{\cal T}_{\rm ideal}^{n}&={\cal D}_{1}\bigg[h^{p}\delta t(-1)^{k+1}\frac{k!(k-1)!}{(2k)!}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\bigg){\cal D}_{p}^{\top}\\ &\hskip 51.6665pt+(-\delta t)^{q+1}\widehat{e}_{\rm RK}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\big(\bm{\alpha}^{n+j}\big)^{q+1}\bigg){\cal D}_{q}^{\top}\bigg].\end{split}$
SM7.3 Numerical results